# Determination of the stably free cancellation property for orders

### v2

@@ 7.5. The reduced norm map
The reduced norm map $\mathrm{nr}:A\rightarrow C$ is defined componentwise (see [Rei03, §9]) and restricts to a group homomorphism $\mathrm{nr}:A^{\times}\rightarrow C^{\times}$. By the Hasse-Schilling-Maas Theorem [CR81, Theorem 7.48] we have $\mathrm{nr}(A^{\times})=C^{\times+}$, where
@@
The last condition means that if $c=(c_{i})$ with $c_{i}\in K_{i}^{\times}$ then $\tau(c_{i})>0$ whenever $i\in\{1,\ldots,r\}$ and $\tau\colon K_{i}\to\mathbb{R}$ is a real embedding such that the corresponding scalar extension $A_{i}\otimes_{K_{i},\tau}\mathbb{R}$ is a full matrix ring over the quaternions. By [CR87, (45.7)] we have that $\mathrm{nr}(\mathcal{M}^{\times})=\mathcal{O}_{C}^{\times}\cap C^{\times+}$.
@@ 7.6. Idèles
Let $\mathfrak{p}$ range over all maximal ideals of $\mathcal{O}$. Let $K_{\mathfrak{p}}$ denote the $\mathfrak{p}$-adic completion of $K$ and let $A_{\mathfrak{p}}=K_{\mathfrak{p}}\otimes_{K}A$. Recall from §2.2 that for an $\mathcal{O}$-lattice $M$, we write $M_{\mathfrak{p}}$ and $\widehat{M}_{\mathfrak{p}}$ for the localisation and completion of $M$ at $\mathfrak{p}$, respectively.
@@
$J(A)=\{(a_{\mathfrak{p}})_{\mathfrak{p}}\in\textstyle{\prod_{\mathfrak{p}}}A^{\times}_{\mathfrak{p}}\mid a_{\mathfrak{p}}\in\widehat{\Lambda}^{\times}_{\mathfrak{p}}\text{ for all but finitely many }\mathfrak{p}\}.$
@@
$A_{\mathfrak{p}}=K_{\mathfrak{p}}\otimes_{K}A$ $\cong\bigoplus_{i=1}^{r}K_{\mathfrak{p}}\otimes_{K}A_{i}\cong\bigoplus_{i=1}^{r}K_{\mathfrak{p}}\otimes_{K}K_{i}\otimes_{K_{i}}A_{i}$
$\cong\bigoplus_{i=1}^{r}\bigoplus_{\mathfrak{P}}\ (K_{i})_{\mathfrak{P}}\otimes_{K_{i}}A_{i}\cong\bigoplus_{i,\mathfrak{P}}A_{i,\mathfrak{P}}$ (7.7)
involving various completions, where, for given $i\in\{1,\ldots,r\}$, $\mathfrak{P}$ runs through all maximal ideals of $\mathcal{O}_{K_{i}}$ dividing $\mathfrak{p}$ and $A_{i,\mathfrak{P}}$ is defined as $(A_{i})_{\mathfrak{P}}$. Using the above isomorphism, we will often write elements of $J(A)$, resp. $A_{\mathfrak{p}}$, as tuples $(a_{i,\mathfrak{P}})_{i,\mathfrak{P}}$, where $\mathfrak{P}$ ranges over all maximal ideals of $\mathcal{O}_{K_{i}}$, resp. over those that contain $\mathfrak{p}$. Similarly, we denote by $J(C)$ the group of idèles of $C$. Again, we have a canonical isomorphism
$C_{\mathfrak{p}}\cong\bigoplus_{i,\mathfrak{P}}\ K_{i,\mathfrak{P}}$
and we will write elements in $J(C)$, resp. $C_{\mathfrak{p}}$, often in the form $(\alpha_{i,\mathfrak{P}})_{i,\mathfrak{P}}$.
By $\mathrm{nr}\colon J(A)\to J(C)$ we also denote the reduced norm map (which translates into the component-wise reduced norm maps $\mathrm{nr}\colon A_{i,\mathfrak{P}}^{\times}\to K_{i,\mathfrak{P}}^{\times}$ after the above identifications). Then by [CR87, (45.8)] we have that $\mathrm{nr}(J(A))=J(C)$.
@@
For $\gamma=(\gamma_{\mathfrak{p}})_{\mathfrak{p}}\in J(A)$ define $\Lambda\gamma=A\cap(\cap_{\mathfrak{p}}\widehat{\Lambda}_{\mathfrak{p}}\gamma_{\mathfrak{p}})$. Then the map $\varepsilon:J(A)\rightarrow\mathrm{Cl}(\Lambda)$ defined by $\alpha\mapsto[\Lambda\alpha]_{\mathrm{st}}$ is a surjective homomorphism. Now let $X$ be a locally free left ideal of $\Lambda$. Then for each $\mathfrak{p}$ there exists $\alpha_{\mathfrak{p}}\in A_{\mathfrak{p}}^{\times}\cap\widehat{\Lambda}_{\mathfrak{p}}$ such $\widehat{X}_{\mathfrak{p}}=\widehat{\Lambda}_{\mathfrak{p}}\alpha_{\mathfrak{p}}$. Set $\alpha=(\alpha_{\mathfrak{p}})_{\mathfrak{p}}$. Then $\Lambda\alpha=X$ and so $\varepsilon(\alpha)=[X]_{\mathrm{st}}$. In fact, $\varepsilon$ factors via $\mathrm{nr}\colon J(A)\to J(C)$ and so under the isomorphism of Theorem 7.4 $(\mathrm{nr}(\alpha_{\mathfrak{p}})_{\mathfrak{p}})\in J(C)$ is a representative of $[X]_{\mathrm{st}}$. Note that we could also have chosen $\alpha_{\mathfrak{p}}\in A^{\times}\cap\Lambda_{\mathfrak{p}}$ such that $X_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}\alpha_{\mathfrak{p}}$. For further details on the claims in this remark, see [CR87, §49A].
@@ 7.7. Review of results of Bley–Boltje
For each $i\in\{1,\ldots,r\}$ we write $\infty_{i}$ for the formal product over real archimedean places $\tau\colon K_{i}\to\mathbb{R}$ such that $\mathbb{R}\otimes_{K_{i},\tau}A_{i}$ is a full matrix ring over the quaternions, and we define the ‘ray modulo $\mathfrak{g}\infty$’ by
@@
$\left(\Lambda/\mathfrak{f}\right)^{\times}\stackrel{{\scriptstyle\nu}}{{\longrightarrow}}I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}\stackrel{{\scriptstyle\mu}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow 0.$ (7.8)
@@
$\tilde{\nu}(x):=(\mathrm{nr}(a)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})\in I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}.$ (7.9)
@@
$\ker(\mu)=\nu((\Lambda/\mathfrak{f})^{\times})=\tilde{\nu}((\Lambda/\mathfrak{f})^{\times}).$ (7.10)
7.8. A condition on $\mathfrak{g}$
For a maximal ideal $\mathfrak{P}$ of $\mathcal{O}_{C}=\mathcal{O}_{K_{1}}\oplus\dotsb\oplus\mathcal{O}_{K_{r}}$ and a finitely generated $\mathcal{O}_{C}$-module $M$ we write $M_{\mathfrak{P}}$ for the localisation of $M$ at $\mathfrak{P}$. Note that for a given $\mathfrak{P}$ there exists a unique $1\leq i\leq r$ and a maximal $\mathcal{O}_{K_{i}}$-ideal $\mathfrak{P}_{i}$ such that
@@ 7.8. A condition on $\mathfrak{g}$
We can and will henceforth identify $M_{\mathfrak{P}}$ with the localisation $(e_{i}M)_{\mathfrak{P}_{i}}$ of the $\mathcal{O}_{K_{i}}$-module $e_{i}M$ at $\mathfrak{P}_{i}$. Note that $M$ is uniquely determined by $M_{\mathfrak{P}}$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ and $M_{\mathfrak{p}}$, where $\mathfrak{p}$ is a maximal ideal of $\mathcal{O}$, is uniquely determined by $M_{\mathfrak{P}}$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ lying above $\mathfrak{p}$.
We shall need the following assumption on $\mathfrak{g}$ in the proof of Theorem 7.10 below.
Assumption (G).
If $\mathfrak{p}$ is a maximal ideal of $\mathcal{O}$ with $\mathfrak{g}_{\mathfrak{p}}\neq\mathcal{O}_{C,\mathfrak{p}}$, then we have $\mathfrak{g}_{\mathfrak{P}}\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ with $\mathfrak{P}\mid\mathfrak{p}$.
@@
Let $L/K$ be an extension of number fields, let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}_{K}$ and let $\mathfrak{P}$ be a prime ideal of $\mathcal{O}_{L}$ lying over $\mathfrak{p}$. We consider the $\mathcal{O}_{K}$-order $\Lambda=\mathcal{O}_{K}+\mathfrak{P}$ and the (unique) maximal $\mathcal{O}_{K}$-order $\mathcal{M}=\mathcal{O}_{L}$ in the $K$-algebra $L$. Then $A=C=L$ and for $\mathfrak{g}=\mathfrak{f}=\mathfrak{P}$ we have $\mathfrak{f}_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}=\mathcal{O}_{C,\mathfrak{p}}$ and assumption (G) is not satisfied whenever there are at least two prime ideals of $\mathcal{O}_{L}$ lying over $\mathfrak{p}$.
@@
We have $\mathfrak{g}_{\mathfrak{p}}=(\mathfrak{f}\cap\mathcal{O}_{C})_{\mathfrak{p}}=\mathfrak{f}_{\mathfrak{p}}\cap\mathcal{O}_{C,\mathfrak{p}}$ since localisations commute with intersections. Thus $1\in\mathfrak{g}_{\mathfrak{p}}$ if and only if $1\in\mathfrak{f}_{\mathfrak{p}}$, which immediately implies (i). Substituting $\mathfrak{P}$ for $\mathfrak{p}$ in this argument shows that $\mathfrak{f}_{\mathfrak{P}}=\mathcal{M}_{\mathfrak{P}}$ if and only if $\mathfrak{g}_{\mathfrak{P}}=\mathcal{O}_{C,\mathfrak{P}}$. By [Rei03, (18.3)], we have that $\mathrm{rad}(\mathcal{M}_{\mathfrak{P}})$ is the unique maximal two-sided ideal of $\mathcal{M}_{\mathfrak{P}}$. Moreover, $\mathfrak{P}$ is clearly the unique maximal ideal of $\mathcal{O}_{C,\mathfrak{P}}$. Together, these last three facts imply (ii). ∎
@@ 7.9. A criterion for stable freeness
Assume that $\mathfrak{g}$ satisfies Assumption (G). Let $X$ be a locally free left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then
@@
where the arrows in the upper triangle are the canonical projections, the isomorphism on the right is that of Theorem 7.4, and the two isomorphisms on the left are those of Theorem 7.6, and $\psi$ is defined to be their composite. Recall that $\mathrm{nr}$, $q_{1}$, $\mu$, and $\varepsilon$ are defined in §7.6, (7.3), (7.8), and Remark 7.5, respectively. The only part of the diagram for which commutativity is not immediately implied by the definitions and Remark 7.5 is the central rectangle, but this follows from the definitions of the isomorphisms as discussed in the proof of [BB06, Theorem 1.5].
@@
Using the canonical isomorphism of (7.7) we write $\alpha_{\mathfrak{p}}=\left(\alpha_{\mathfrak{P}}\right)_{\mathfrak{P}\mid\mathfrak{p}}$ with $\alpha_{\mathfrak{P}}\in A_{\mathfrak{P}}$ for the local basis elements $\alpha_{\mathfrak{p}}$ chosen as above. We will apply the above recipe for $\gamma_{\mathfrak{P}}=\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}})$. The result of [BB06, Corollary 2.4] implies that for $\mathfrak{P}\mid\mathfrak{g}$, we have $\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}})\equiv 1\bmod{\mathfrak{g}_{\mathfrak{P}}}$, so we can and do choose $\xi=1$.
@@
${v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}}))}=v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\beta))$ (7.11)
@@
Hence there exists an element $\gamma\in\mathcal{M}_{\mathfrak{p}}$ such that $\gamma\beta\equiv 1\bmod{\mathfrak{f}_{\mathfrak{p}}}$. It follows that $\gamma\beta\equiv 1\bmod{\mathfrak{f}_{\mathfrak{P}}}$ for all $\mathfrak{P}\mid\mathfrak{p}$. From [BB06, Corollary 2.4] we obtain $\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\gamma\beta)\equiv 1\bmod{\mathfrak{g}_{\mathfrak{P}}}$ for all $\mathfrak{P}\mid\mathfrak{p}$. Since $\mathfrak{g}_{\mathfrak{p}}\neq\mathcal{O}_{C,\mathfrak{p}}$ by Lemma 7.9, assumption (G) implies that $\mathfrak{g}_{\mathfrak{P}}\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$. Hence $v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\beta))=0$. Since $\alpha_{\mathfrak{p}}\equiv 1\bmod{\mathfrak{f}_{\mathfrak{p}}}$ the same argument shows that $v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}}))=0$ for all maximal ideals $\mathfrak{P}\mid\mathfrak{p}$.

### v3

@@ 7.5. The reduced norm map
The reduced norm map $\mathrm{nr}:A\rightarrow C$ is defined componentwise (see [Rei03, §9]) and restricts to a group homomorphism $\mathrm{nr}:A^{\times}\rightarrow C^{\times}$. By the Hasse–Schilling–Maass Theorem [CR81, Theorem 7.48] we have $\mathrm{nr}(A^{\times})=C^{\times+}$, where
@@
The last condition means that if $c=(c_{i})$ with $c_{i}\in K_{i}^{\times}$ then $\tau(c_{i})>0$ whenever $i\in\{1,\ldots,r\}$ and $\tau\colon K_{i}\to\mathbb{R}$ is a real embedding such that the corresponding scalar extension $A_{i}\otimes_{K_{i},\tau}\mathbb{R}$ is isomorphic to a full matrix algebra over $\mathbb{H}$. By [CR87, (45.7)] we have that $\mathrm{nr}(\mathcal{M}^{\times})=\mathcal{O}_{C}^{\times}\cap C^{\times+}$.
@@ 7.6. Idèles
Let $\mathfrak{p}$ range over all maximal ideals of $\mathcal{O}$. Recall from §2.2 that for an $\mathcal{O}$-lattice $M$, we write $M_{\mathfrak{p}}$ and $\widehat{M}_{\mathfrak{p}}$ for the localization and completion of $M$ at $\mathfrak{p}$, respectively. Extending this notation, let $\widehat{K}_{\mathfrak{p}}$ denote the $\mathfrak{p}$-adic completion of $K$ and let $\widehat{A}_{\mathfrak{p}}=\widehat{K}_{\mathfrak{p}}\otimes_{K}A$.
@@
$J(A)=\{(a_{\mathfrak{p}})_{\mathfrak{p}}\in\textstyle{\prod_{\mathfrak{p}}}\widehat{A}^{\times}_{\mathfrak{p}}\mid a_{\mathfrak{p}}\in\widehat{\Lambda}^{\times}_{\mathfrak{p}}\text{ for all but finitely many }\mathfrak{p}\}.$
@@
$\widehat{A}_{\mathfrak{p}}=\widehat{K}_{\mathfrak{p}}\otimes_{K}A$ $\cong\bigoplus_{i=1}^{r}\widehat{K}_{\mathfrak{p}}\otimes_{K}A_{i}\cong\bigoplus_{i=1}^{r}\widehat{K}_{\mathfrak{p}}\otimes_{K}K_{i}\otimes_{K_{i}}A$
(7.4) $\cong\bigoplus_{i=1}^{r}\bigoplus_{\mathfrak{P}}\ \widehat{(K_{i})}_{\mathfrak{P}}\otimes_{K_{i}}A_{i}\cong\bigoplus_{i,\mathfrak{P}}\widehat{A}_{i,\mathfrak{P}}$
involving various completions, where, for given $i\in\{1,\ldots,r\}$, $\mathfrak{P}$ runs through all maximal ideals of $\mathcal{O}_{K_{i}}$ dividing $\mathfrak{p}$ and $\widehat{A}_{i,\mathfrak{P}}$ is defined as $\widehat{(A_{i})}_{\mathfrak{P}}$. Using (7.4), we will often write elements of $J(A)$, resp. $\widehat{A}_{\mathfrak{p}}$, as tuples $(a_{i,\mathfrak{P}})_{i,\mathfrak{P}}$, where $\mathfrak{P}$ ranges over all maximal ideals of $\mathcal{O}_{K_{i}}$, resp. over those that contain $\mathfrak{p}$. Similarly, we denote by $J(C)$ the group of idèles of $C$. Again, we have a canonical isomorphism
$\widehat{C}_{\mathfrak{p}}\cong\bigoplus_{i,\mathfrak{P}}\ \widehat{K}_{i,\mathfrak{P}}$
and we will write elements in $J(C)$, resp. $\widehat{C}_{\mathfrak{p}}$, often in the form $(\alpha_{i,\mathfrak{P}})_{i,\mathfrak{P}}$.
By $\mathrm{nr}\colon J(A)\to J(C)$ we also denote the reduced norm map (which translates into the componentwise reduced norm maps $\mathrm{nr}\colon\widehat{A}_{i,\mathfrak{P}}^{\times}\to\widehat{K}_{i,\mathfrak{P}}^{\times}$ after the above identifications). Then by [CR87, (45.8)] we have that $\mathrm{nr}(J(A))=J(C)$.
@@
For $\gamma=(\gamma_{\mathfrak{p}})_{\mathfrak{p}}\in J(A)$ define $\Lambda\gamma=A\cap(\cap_{\mathfrak{p}}\widehat{\Lambda}_{\mathfrak{p}}\gamma_{\mathfrak{p}})$. Then the map $\varepsilon:J(A)\rightarrow\mathrm{Cl}(\Lambda)$ defined by $\alpha\mapsto[\Lambda\alpha]_{\mathrm{st}}$ is a surjective homomorphism. Now let $X$ be a locally free left ideal of $\Lambda$. Then for each $\mathfrak{p}$ there exists $\alpha_{\mathfrak{p}}\in\widehat{A}_{\mathfrak{p}}^{\times}\cap\widehat{\Lambda}_{\mathfrak{p}}$ such $\widehat{X}_{\mathfrak{p}}=\widehat{\Lambda}_{\mathfrak{p}}\alpha_{\mathfrak{p}}$. Set $\alpha=(\alpha_{\mathfrak{p}})_{\mathfrak{p}}$. Then $\Lambda\alpha=X$ and so $\varepsilon(\alpha)=[X]_{\mathrm{st}}$. In fact, $\varepsilon$ factors via $\mathrm{nr}\colon J(A)\to J(C)$ and so under the isomorphism of Theorem 7.4 $(\mathrm{nr}(\alpha_{\mathfrak{p}})_{\mathfrak{p}})\in J(C)$ is a representative of $X]_{\mathrm{st}}$. Note that we could also have chosen $\alpha_{\mathfrak{p}}\in A^{\times}\cap\Lambda_{\mathfrak{p}}$ such that $X_{\mathfrak{p}}=\Lambda_{\mathfrak{p}}\alpha_{\mathfrak{p}}$. For further details on the claims in this remark, see [CR87, §49A].
@@ 7.7. Review of results of Bley–Boltje
For each $i\in\{1,\ldots,r\}$ we write $\infty_{i}$ for the formal product over real archimedean places $\tau\colon K_{i}\to\mathbb{R}$ such that $\mathbb{R}\otimes_{K_{i},\tau}A_{i}$ is a full matrix algebra over $\mathbb{H}$, and we define the ‘ray modulo $\mathfrak{g}\infty$’ by
@@
(7.5) $\left(\Lambda/\mathfrak{f}\right)^{\times}\stackrel{{\scriptstyle\nu}}{{\longrightarrow}}I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}\stackrel{{\scriptstyle\mu}}{{\longrightarrow}}\mathrm{Cl}(\Lambda)\longrightarrow 0.$
@@
(7.6) $\tilde{\nu}(x):=(\mathrm{nr}(a)\mathcal{O}_{C}\bmod P_{\mathfrak{g}}^{+})\in I_{\mathfrak{g}}/P_{\mathfrak{g}}^{+}.$
@@
(7.7) $\ker(\mu)=\nu((\Lambda/\mathfrak{f})^{\times})=\tilde{\nu}((\Lambda/\mathfrak{f})^{\times}).$
7.8. A technical condition on $\mathfrak{g}$
For a maximal ideal $\mathfrak{P}$ of $\mathcal{O}_{C}=\mathcal{O}_{K_{1}}\oplus\dotsb\oplus\mathcal{O}_{K_{r}}$ and a finitely generated $\mathcal{O}_{C}$-module $M$, we write $M_{\mathfrak{P}}$ for the localization of $M$ at $\mathfrak{P}$. Note that for a given $\mathfrak{P}$ there exists a unique $1\leq i\leq r$ and a maximal $\mathcal{O}_{K_{i}}$-ideal $\mathfrak{P}_{i}$ such that
@@ 7.8. A technical condition on $\mathfrak{g}$
We can and will henceforth identify $M_{\mathfrak{P}}$ with the localization $(e_{i}M)_{\mathfrak{P}_{i}}$ of the $\mathcal{O}_{K_{i}}$-module $e_{i}M$ at $\mathfrak{P}_{i}$. Note that $M$ is uniquely determined by $M_{\mathfrak{P}}$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ and that $M_{\mathfrak{p}}$, where $\mathfrak{p}$ is a maximal ideal of $\mathcal{O}$, is uniquely determined by $M_{\mathfrak{P}}$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ lying above $\mathfrak{p}$.
We make the following non-standard definition. It may be viewed as a technical assumption on $\mathfrak{g}$ that is needed in the proof of Theorem 7.11 below. In algorithmic applications, $\mathfrak{g}$ will be adjusted to ensure that it satisfies this assumption.
Definition 7.8.
The ideal $\mathfrak{g}$ is said to be locally radical if for each maximal ideal $\mathfrak{p}$ of $\mathcal{O}$ with $\mathfrak{g}_{\mathfrak{p}}\neq\mathcal{O}_{C,\mathfrak{p}}$, we have $\mathfrak{g}_{\mathfrak{P}}\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$ for all maximal ideals $\mathfrak{P}$ of $\mathcal{O}_{C}$ with $\mathfrak{P}\mid\mathfrak{p}$.
@@
Let $L/K$ be an extension of number fields, let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}_{K}$ and let $\mathfrak{P}$ be a prime ideal of $\mathcal{O}_{L}$ lying over $\mathfrak{p}$. We consider the $\mathcal{O}_{K}$-order $\Lambda=\mathcal{O}_{K}+\mathfrak{P}$ and the (unique) maximal $\mathcal{O}_{K}$-order $\mathcal{M}=\mathcal{O}_{L}$ in the $K$-algebra $L$. Then $A=C=L$ and for $\mathfrak{g}=\mathfrak{f}=\mathfrak{P}$ we have $\mathfrak{f}_{\mathfrak{p}}\neq\mathcal{M}_{\mathfrak{p}}=\mathcal{O}_{C,\mathfrak{p}}$ and $\mathfrak{g}$ is locally radical if and only if there is a unique prime ideal of $\mathcal{O}_{L}$ lying over $\mathfrak{p}$.
@@
We have $\mathfrak{g}_{\mathfrak{p}}=(\mathfrak{f}\cap\mathcal{O}_{C})_{\mathfrak{p}}=\mathfrak{f}_{\mathfrak{p}}\cap\mathcal{O}_{C,\mathfrak{p}}$ since localizations commute with finite intersections. Thus $1\in\mathfrak{g}_{\mathfrak{p}}$ if and only if $1\in\mathfrak{f}_{\mathfrak{p}}$, which immediately implies (i). Substituting $\mathfrak{P}$ for $\mathfrak{p}$ in this argument shows that $\mathfrak{f}_{\mathfrak{P}}=\mathcal{M}_{\mathfrak{P}}$ if and only if $\mathfrak{g}_{\mathfrak{P}}=\mathcal{O}_{C,\mathfrak{P}}$. By [Rei03, (18.3)], we have that $\mathrm{rad}(\mathcal{M}_{\mathfrak{P}})$ is the unique maximal two-sided ideal of $\mathcal{M}_{\mathfrak{P}}$. Moreover, $\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$ is the unique maximal ideal of $\mathcal{O}_{C,\mathfrak{P}}$. Together, these last three facts imply (ii). ∎
@@ 7.9. A criterion for stable freeness
Assume that $\mathfrak{g}$ is locally radical. Let $X$ be a locally free left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then
@@
where the arrows in the upper triangle are the canonical projections, the isomorphism on the right is that of Theorem 7.4, the two isomorphisms on the left are those of Theorem 7.6, and $\psi$ is defined to be their composite. Recall that $\mathrm{nr}$, $q_{1}$, $\mu$, and $\varepsilon$ are defined in §7.6, (7.2), (7.5), and Remark 7.5, respectively. The only part of the diagram for which commutativity is not immediately implied by the definitions and Remark 7.5 is the central rectangle, but this follows from the definitions of the isomorphisms as discussed in the proof of [BB06, Theorem 1.5].
@@
Using the canonical isomorphism of (7.4) we write $\alpha_{\mathfrak{p}}=\left(\alpha_{\mathfrak{P}}\right)_{\mathfrak{P}\mid\mathfrak{p}}$ with $\alpha_{\mathfrak{P}}\in\widehat{A}_{\mathfrak{P}}$ for the local basis elements $\alpha_{\mathfrak{p}}$ chosen as above. We will apply the above recipe for $\gamma_{\mathfrak{P}}=\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}})$. The result of [BB06, Corollary 2.4] implies that for $\mathfrak{P}\mid\mathfrak{g}$, we have $\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}})\equiv 1\bmod{\mathfrak{g}_{\mathfrak{P}}}$, so we can and do choose $\xi=1$.
@@
(7.8) ${v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}}))}=v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\beta))$
@@
Hence there exists an element $\gamma\in\mathcal{M}_{\mathfrak{p}}$ such that $\gamma\beta\equiv 1\bmod{\mathfrak{f}_{\mathfrak{p}}}$. It follows that $\gamma\beta\equiv 1\bmod{\mathfrak{f}_{\mathfrak{P}}}$ for all $\mathfrak{P}\mid\mathfrak{p}$. From [BB06, Corollary 2.4] we obtain $\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\gamma\beta)\equiv 1\bmod{\mathfrak{g}_{\mathfrak{P}}}$ for all $\mathfrak{P}\mid\mathfrak{p}$. Since $\mathfrak{g}_{\mathfrak{p}}\neq\mathcal{O}_{C,\mathfrak{p}}$ by Lemma 7.10, the assumption that $\mathfrak{g}$ is locally radical implies that $\mathfrak{g}_{\mathfrak{P}}\subseteq\mathfrak{P}\mathcal{O}_{C,\mathfrak{P}}$. Hence $v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\beta))=0$. Since $\alpha_{\mathfrak{p}}\equiv 1\bmod{\mathfrak{f}_{\mathfrak{p}}}$ the same argument shows that $v_{\mathfrak{P}}(\mathrm{nr}_{A_{\mathfrak{P}}/L_{\mathfrak{P}}}(\alpha_{\mathfrak{P}}))=0$ for all maximal ideals $\mathfrak{P}\mid\mathfrak{p}$.